OpenStudy (anonymous):

What value completes the square for the following? x2 + 8x + __ A. 4 B. 8 C. 16 D. 32

OpenStudy (atlas):

Can you tell what is (a+b)^2 = ?

OpenStudy (anonymous):

no

OpenStudy (atlas):

can you multiply (a+b)(a+b)

OpenStudy (anonymous):

yes

OpenStudy (atlas):

what do you get?

OpenStudy (anonymous):

well i dont kow which ones are a and b

OpenStudy (anonymous):

It is 16 x^2+8x+16=(x+4)^2